OpenStudy (dls):

Test whether the following function is odd/even/either/neither \LARGE \frac{\sin^3x}{[\frac {x}{\pi}]+\frac{1}{2}}

OpenStudy (dls):

@ParthKohli @AravindG

OpenStudy (anonymous):

check putting -x instead of x, if the fucntion changes, then it's odd if not then even

OpenStudy (dls):

it is not that easy

OpenStudy (aravindg):

sorry keyboard nt working

OpenStudy (dls):

:/

OpenStudy (anonymous):

sin(-x)=-sin(x)

OpenStudy (dls):

\[\LARGE f(-x)=\frac{-\sin^3x}{[\frac {-x}{\pi}]+[\frac{1}{2}]}\]

OpenStudy (anonymous):

so... it is?

OpenStudy (dls):

u say?

OpenStudy (anonymous):

if f(x)=f(-x) it's even

OpenStudy (dls):

how is it =fx?

OpenStudy (anonymous):

it is odd

OpenStudy (dls):

:/

OpenStudy (anonymous):

f(x) refers to the "function of x"

OpenStudy (dls):

yes

OpenStudy (anonymous):

your result is \[f(x)neqf(-x)\]

OpenStudy (anonymous):

not equal

OpenStudy (anonymous):

if it would be equal then even

OpenStudy (dls):

@terenzreignz

OpenStudy (aravindg):